Find the gravitational field on the surface of an alien planet that has an estimated mass of 8.5 x 10^26 kg and radius of 3.5 x 10^4 km. Show all work.
Answer #1

o(r) = Gm,

= (6.67 * 10-11) * (8.5 * 1026) / (3.5*107)2 = 4.62 * 10 = 46.2 N/kg